OpenStudy (anonymous):

Form the equation of the line with the following conditions: passes through (3,2) and (1,-4) ?

OpenStudy (anonymous):

Alright first you must find the slope of the line using that last equation you used

OpenStudy (saifoo.khan):

slope!

OpenStudy (anonymous):

yo yo yo

OpenStudy (anonymous):

m= (y1-y2) over (x1-x2) right?

OpenStudy (saifoo.khan):

right

OpenStudy (anonymous):

You should get 3 = m

OpenStudy (anonymous):

I did. I'm clueless as to what to do after.

OpenStudy (anonymous):

It says form the equation? I'm sorry, I missed tons of classes and am only cramming now.:)

OpenStudy (anonymous):

alright so there is another equation you need- Point slope form which is \[y-y _{1}=m(x-x _{1)}\]

OpenStudy (anonymous):

have you seen this equation?

OpenStudy (anonymous):

I have. Just substitution, right?

OpenStudy (anonymous):

yes, with either point. With the first, you should get y-2=3(x-3) y-2=3x-9 y=3x-7

OpenStudy (anonymous):

I did. Thank you so much for you help! Salamat! :)

OpenStudy (anonymous):

*your

OpenStudy (anonymous):

with the second you'll get y+4=3(x-1) y+4=3x-3 y=3x-7 *just showing either way

OpenStudy (anonymous):

Thank you, thank you, thank you! :)